### v1

# OccLE: Label-Efficient 3D Semantic Occupancy Prediction

Abstract
3D semantic occupancy prediction offers an intuitive and efficient scene understanding and has attracted significant interest in autonomous driving perception. Existing approaches either rely on full supervision, which demands costly voxel-level annotations, or on self-supervision, which provides limited guidance and yields suboptimal performance. To address these challenges, we propose OccLE, a Label-Efficient 3D Semantic Occupancy Prediction that takes images and LiDAR as inputs and maintains high performance with limited voxel annotations. Our intuition is to decouple the semantic and geometric learning tasks and then fuse the learned feature grids from both tasks for the final semantic occupancy prediction. Therefore, the semantic branch distills 2D foundation model to provide aligned pseudo labels for 2D and 3D semantic learning. The geometric branch integrates image and LiDAR inputs in cross-plane synergy based on their inherency, employing semi-supervision to enhance geometry learning. We fuse semantic-geometric feature grids through Dual Mamba and incorporate a scatter-accumulated projection to supervise unannotated prediction with aligned pseudo labels. Experiments show that OccLE achieves competitive performance with only 10% of voxel annotations, reaching a mIoU of 16.59% on the SemanticKITTI validation set.
Figure 1: Label-efficient 3D semantic occupancy prediction aims to achieve high performance using limited voxel annotations and aligned pseudo label. We propose OccLE, a novel learning paradigm that decouples semantic and geometric learning and fuse their feature grids for the final prediction.
Keywords Autonomous Driving  $\cdot$ 3D Semantic Occupancy Prediction  $\cdot$ Label-Efficient Learning
1 Introduction
3D perception task is foundational cornerstone for autonomous driving systems. Among various perception methods, 3D semantic occupancy prediction [1, 2, 3] has garnered significant attention for providing intuitive and efficient scene understanding for downstream tasks. This task aims to estimate the occupancy status and semantic label of each voxel in a 3D grid, given input from 2D images, 3D LiDAR, or a combination of both.
Previous studies [4, 5] proposed numerous supervised learning paradigms with diverse modal inputs. A notable challenge is that their high performance relies heavily on extensive voxel annotations, which are both costly and labor-intensive. The cubic complexity of the voxel grid leads to a substantial workload during manual annotation. Even with pre-annotated 3D labels generated using auto-labeling assistants, one hour of human labor just only annotates 10 frames [6, 7]. This issue limits the scalability of supervised methods for real-world deployment, especially in terms of robustness.
Recent advancements in 3D semantic occupancy prediction have shifted towards a label-free paradigm. These self-supervised methods leverage existing vision foundation models to generate various pseudo labels for auxiliary supervision, such as image semantic segmentation [8], depth information [9], and LiDAR segmentation [10]. To bridge the 3D representation with 2D pseudo label, specific 3D representation formats have been proposed to facilitate volume rendering, including signed distance functions[7] and 3D Gaussians [11]. The core strategy of these self-supervised methods is to supplement pseudo semantic supervision. They project LiDAR segmentation annotations from the same dataset into 2D pseudo labels [12, 6]. However, this pseudo labels generation still is expensive and provides only sparse supervision. They also employ a 2D vision foundation model to generate 2D pseudo labels [7], which suffers from class misalignment. For the overall pipeline, they supervise only the final prediction using 2D pseudo labels and do not apply separate supervision to geometric and semantic feature learning. Thus, this non-decoupled design leads to insufficient geometric learning, which reduces volume-rendering quality and weakens semantic supervision.
To address these challenges, we propose OccLE, a label-efficient 3D semantic occupancy prediction method, as illustrated in Fig. 1, which uses limited voxel annotations while maintaining high performance. We design the label-efficient learning paradigm: taking image and LiDAR scan inputs, fully decoupling semantic and geometric feature learning, fusing their feature grids for 3D semantic occupancy prediction, and supervising with limited voxel annotations and aligned pseudo labels. Specifically, OccLE consists of the semantic branch, the geometric branch, and semantic-geometric feature grid fusion. The semantic branch distills dataset-specific and open-vocabulary 2D foundation models to produce aligned pseudo labels for supervising 2D and 3D semantic learning, independently of geometry. The geometric branch integrates image and LiDAR and inputs in cross-plane synergy based on their inherent characteristics, employing semi-supervision to enhance geometry learning. The semantic-geometric feature grid fusion employs Dual Mamba to fuse the two feature grids for lightweight long-range relationship modeling, and introduces a scatter-accumulated projection to guide unannotated predictions with aligned pseudo labels, without relying on any specific feature grid format. In summary, the main contributions of this work are summarized as follows: In summary, the main contributions of this work are summarized as follows:
• We introduce a label-efficient 3D semantic occupancy prediction task and a learning paradigm that decouples semantic and geometric learning, followed by their synergistic integration. The model is supervised using limited voxel annotations and aligned pseudo labels to achieve high performance.
• We present three techniques: distilling 2D foundation model for 2D and 3D semantic learning; semi-supervised geometry learning using cross-plane image and LiDAR feature synergy; and semantic-geometric feature grids fusion with auxiliary supervision by scatter-accumulated projection.
• Experimental results show that OccLE achieves competitive performance compared to fully supervised methods, attaining 16.59 % mIoU on the SemanticKITTI validation set using only 10 % of the voxel annotations.
2 Related Works
2.1 3D Semantic Occupancy Prediction
3D semantic occupancy prediction, also referred to as the semantic scene completion (SSC), was first introduced by SSCNet [1]. In SSCNet, image features are lifted into a 3D representation to predict voxel-level semantics. With the advancement of autonomous driving [2, 3], SSC has attracted significant attention for providing clear and efficient scene representations. Based on the type of input, SSC methods can be categorized into camera-based and multi-modal methods.
For camera-based 3D semantic occupancy prediction, a key challenge lies in lifting 2D to 3D representations. Previous works [13, 14, 15] utilize calibration and depth information to establish explicit backwards projections. ViewFormer [16] proposes a learning-first attention mechanism to aggregate multi-view features. Some studies [17, 18] represent scenes using sparse 3D semantic Gaussians. TPVFormer [19] introduces three perpendicular views (TPV) to enhance the interaction of image features.
Since dimensional transformation based on depth estimation is ill-posed for camera-based methods, multi-modal method incorporates LiDAR and radar inputs to obtain accurate distance measurements. Its essential research is modal fusion. OccFusion [5] directly concatenates feature channels. MetaOcc [20] learn global and local alignment to enhance fused representations. Some studies [21, 22] advance fusion by employing point-to-point queries from 3D to 2D. OccGen [23] employs a diffusion model to refine the fused features.
Building on prior work, OccLE takes camera images and a LiDAR scans as inputs, applies backward projection to derive 3D feature grids from 2D image feature maps, and proposes a lightweight yet effective fusion strategy in the geometry learning termed cross-plane image and LiDAR feature synergy.
2.2 Label-efficient learning
Label-efficient learning seeks to minimize the reliance on large amounts of label while maintaining high model performance. Prior works have proposed novel learning paradigms to achieve this, such as semi-supervised learning and self-supervised learning. Semi-supervised learning [24] combines a small labeled dataset with abundant unlabeled data to improve model generalization, while self-supervised learning [25] generates supervision directly from unlabeled data through pretext tasks.
In label-efficient learning for 3D semantic occupancy prediction, self-supervision methods render 3D feature into 2D representations, and supervise with labels from other tasks. To facilitate rendering, SelfOcc [7] introduces a signed distance function-based representation, while GaussTR [11] proposes a 3D Gaussian-based representation. For auxiliary supervision, OccFlowNet [12] utilizes projected semantic point clouds, whereas RenderOcc [6] relies on depth estimation and image segmentation.
Inspired by the above, OccLE utilizes the pseudo label from image segmentation as a key supervision to enable label-efficient learning. We distill 2D foundation model for multi-modal semantic learning, while supervising the accumulated projection in the final prediction.
3 Methodology
3.1 Overview
Figure 2: The overview of OccLE. First, we distill 2D foundation models to predict aligned pseudo labels for supervising 2D and 3D semantic learning. Next, we propose cross-plane image and LiDAR feature synergy and apply semi-supervision to learn geometry. Finally, we fuse semantic and geometric feature grids via Dual Mamba and supervise the unanotated prediction with aligned pseudo label using scatter-accumulated projection.
We aim to take multi-frame camera images ${\left\{{{{\bf{I}}_{t-i}}}\right\}_{i=0}^{N-1}\in{{\mathbb{R}}^{h\times w}}}$, multi-frame point clouds ${\left\{{{{\bf{P}}_{t-i}}}\right\}_{i=0}^{N-1}\in{{\mathbb{R}}^{n\times 4}}}$, camera calibration ${\left\{{{\bf{K}}{\rm{,}}{\bf{T}}}\right\}}$ as inputs to predict the 3D semantic occupancy ${{{\bf{Y}}_{t}}\in{\left\{{{c_{0}},{c_{1}},\cdots,{c_{M-1}}}\right\}^{H\times W\times Z}}}$ at timestep ${t}$. To enable label-efficient training, we supervise the model using only limited voxel annotations (e.g., 10% ${\left\{{{{{\bf{\bar{Y}}}}_{t}}}\right\}}$ of all samples). Here, ${{\bf{K}}{\rm{,}}{\bf{T}}}$ represent the intrinsic and extrinsic camera matrices, respectively; ${\left\{{{{{\bf{\bar{Y}}}}_{t}}}\right\}}$ is the ground truth of 3D semantic occupancy prediction. ${N}$ denotes frame number; ${M}$ denotes class number, ${h}$ and ${w}$ denote the height and width of camera image; ${H}$, ${W}$, and ${Z}$ represent the length, width, height of voxel grid. Our method is formulated as ${{{\bf{Y}}_{t}}=\Theta\left({\left\{{{{\bf{I}}_{t-i}}}\right\}_{i=0}^{N-1},\left\{{{{\bf{P}}_{t-i}}}\right\}_{i=0}^{N-1},{\bf{K}}{\rm{,}}{\bf{T}}}\right)}$, where ${\Theta}$ is our proposed OccLE. For clarity, we omit the timestep subscript ${t}$ for all variables in the following description.
The overview of OccLE is illustrated in Fig. 2. For simplicity, only a single input frame is shown. To enable label-efficient learning, our pipeline comprises three components: a semantic branch, a geometric branch, and a semantic-geometric feature grid fusion module. We decouple the semantic and geometric learning with distinct supervision strategies, and fuse their feature grids to predict 3D semantic occupancy. In the semantic branch, we distill 2D foundation models to predict aligned pseudo labels and supervise 2D and 3D semantic feature learning. In the geometric branch, we synergize image and LiDAR features based on their inherency and employ semi-supervision to strengthen geometry learning. In the semantic-geometric feature grid fusion module, we employ Dual Mamba to fuse feature grids and employ scatter-accumulated projection to supervise unanotated predictions with aligned pseudo labels. The semantic branch, the geometric branch, and the semantic-geometric feature grid fusion will be discussed in Sec. 3.2, Sec. 3.3, and Sec. 3.4, respectively.
3.2 Distill 2D Foundation Models for 2D and 3D Semantic Learning
3D semantic occupancy prediction is implemented through the entanglement of geometric and semantic features [4]. The geometric feature indicates whether a voxel is occupied, while the semantic feature specifies the class occupying the voxel. 3D semantic occupancy prediction maps the feature grid to voxel semantic labels. Achieving high accuracy traditionally requires extensive voxel annotations to guide the mapping, but obtaining such annotations is costly. To mitigate this requirement, we notice that the image semantic segmentation maps the feature map to pixel semantic label and benefits from large annotated datasets. Therefore, we aim to distill 2D foundation model for 2D and 3D semantic learning, where the 2D foundation model predicts aligned pseudo labels to supervise the semantic learning of image and LiDAR scan inputs.
2D foundation models may not share the same class definitions as 3D semantic occupancy prediction. For example, a 2D foundation model trained on the Cityscapes dataset [26] does not include categories such as bicyclist or truck when applied to the SemanticKITTI dataset [27]. To resolve this class unalignment, we integrate both dataset-specific and open-vocabulary 2D foundation models. The dataset-specific one produces primary labels for shared classes, while the open-vocabulary one produces auxiliary labels for unaligned classes. Specifically, we utilize MSeg [28] as a dataset-specific model and SAM2 [8] for open-vocabulary model. We adapt MSeg universal class classification to the task-specific classes, input the unaligned class name prompts into SAM2, and merge outputs using Boolean operations to generate aligned pseudo labels ${{\bf{\bar{s}}}}$.
We employ ${{\bf{\bar{s}}}}$ to fully supervise semantic learning of images and LiDAR scans, enabling lightweight deployment via distillation. Specifically, we employ a UNet [29] ${{{\cal G}_{s2d}}}$ to extract features map ${{\bf{F}}_{I}^{s}}$ from the image ${{\bf{I}}}$. We adopt the voxelization in VoxelNet [30] and utilize a sparse UNet3D [31] ${{{\cal G}_{s3d}}}$ to extract features grid ${{\bf{F}}_{P}^{s}}$ from the LiDAR scan ${{\bf{P}}}$. The mathematical expression is given by:
[…]
where ${{\cal V}}$ refers to the voxelization process, which randomly samples up to 35 points per voxel. Each point includes its position, intensity, and the offset from the voxel center.
The semantic branch predicts image segmentation ${{\bf{s}}}$ and LiDAR segmentation ${{\bf{S}}}$ using a 2D header and 3D header. To supervise ${{\bf{S}}}$, we lift ${{\bf{\bar{s}}}}$ to ${{\bf{\bar{S}}}}$ as follows:
${{\bf{\bar{S}}}}={\cal F}\left({{\bf{\bar{s}}},{\bf{p}}}\right),{\bf{p}}={\bf{K}}\left({{\bf{T}}\cdot{\bf{P}}}\right)$ (2)
where ${{\cal F}}$ denotes the sampling function, ${{\bf{P}}}$ denote the voxel grid coordinates, ${{\bf{p}}}$ represent the corresponding projected coordinates on the image plane.
3.3 Semi-Supervised Geometry Learning
Prior works [4, 32] depend on absolute depth in learning geometry for 3D semantic occupancy prediction. Because depth estimation is ill-posed, reducing voxel annotations exacerbates uncertainty in this sequential prediction. Since depth is sourced from the LiDAR scan projected on the image plane, we directly use image and LiDAR scan inputs, propose a lightweight and modality-complementary geometry learning module, termed cross-plane image and LiDAR feature synergy, and employ a semi-supervision to learn an accurate geometric feature grid with limited voxel annotations.
3.3.1 Cross-Plane Image and LiDAR Feature Synergy
First, we examine the characteristics of the image and LiDAR scan inputs. As shown by the red boxes in Fig. 3a and Fig. 3b, the frontal-view LiDAR scan is sparse and omits information in the upper region, whereas the image offers continuous context. In the bird’s-eye view (BEV), the image cannot resolve distances along each ray, while the LiDAR scan provides accurate range measurements. In summary, the image is reliable on the ${yz}$-plane and the LiDAR scan is reliable along the ${x}$-axis in world coordinates. Motivated by their inherency and inspired by [33], we project each their feature grid into its optimal plane in the TPV and fuse them via multiplicative integration.
Figure 3: Illustration of geometry learning. (a) Frontal view feature comparison. (b) BEV view feature comparison. (c) The cross-plane image and LiDAR feature synergy.
The cross-plane image and LiDAR feature synergy is illustrated in Fig. 3c. We voxelize multi-frame LiDAR scans using the VoxelNet [30] and process them with a sparse 3D encoder to obtain LiDAR feature grids. In parallel, we extract feature maps from multi-frame images using a 2D encoder and lift into the feature grid defined in Equ. 2. We fuse the multi-frame feature grids into a single grid via a Conv1D layer. For the LiDAR feature grid, we stack along the ${y}$ and ${z}$ axes and apply a Conv2D layer to produce feature maps ${{\bf{F}}_{I,xz}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{Z_{1}}}}}$ and ${{\bf{F}}_{I,xy}^{g}\in{{\mathbb{R}}^{{W_{1}}\times{Z_{1}}}}}$. For the image feature grid, we stack along the ${x}$-axis and apply a Conv2D layer to generate feature map ${{\bf{F}}_{P,yz}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}}$. To enhance ${{\bf{F}}_{P,xy}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}}$, we apply self-attention (SA) along the ${y}$ and ${z}$ axes, yielding feature maps ${{\bf{F}}_{P,yzy}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}}$, ${{\bf{F}}_{P,yzz}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}}$ as follows:
$\left\{\begin{array}[]{l}{\bf{F}}_{P,yzy}^{g}={\bf{F}}_{P,yz}^{g}+SA\left({{\bf{F}}_{P,yz}^{g}}\right)\\ {\bf{F}}_{P,yzz}^{g}={\bf{F}}_{P,yz}^{g}+SA{\left({{\bf{F}}{{{}_{P,yz}^{g}}^{T}}}\right)^{T}}\end{array}\right.$ (3)
We fuse the image and LiDAR feature maps via matrix multiplication in the TPV. The fused feature grids ${{\bf{F}}_{{f_{1}}}^{g},{\bf{F}}_{{f_{2}}}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}\times{Z_{1}}}}}$ are defined as follows:
$\left\{\begin{array}[]{l}{\bf{F}}_{{f_{1}}}^{g}={\bf{F}}_{I,xz}^{g}\times{\bf{F}}_{P,yzz}^{g}\\ {\bf{F}}_{{f_{2}}}^{g}={\bf{F}}_{I,xy}^{g}\times{\bf{F}}_{P,yzy}^{g}\end{array}\right.$ (4)
We add ${{\bf{F}}_{{f_{1}}}^{g}}$ and ${{\bf{F}}_{{f_{2}}}^{g}}$ to generate the geometric feature grid ${{\bf{F}}_{{f}}^{g}}$ and predict geometry ${{\bf{G}}\in{{\mathbb{R}}^{H\times W\times Z}}}$ via a 3D header.
In the proposed workflow, feature grids are converted into feature maps by stacking and subsequently reconstructed into feature grids during the final fusion. In contrast to the 3D attention mechanisms adopted in [4, 32, 34], the cross-plane image and LiDAR feature synergy employs only Conv1D, Conv2D, and SA (with computational complexity ${{\cal O}\left({{H_{1}}^{2}}\right)}$ or ${{\cal O}\left({{W_{1}}^{2}}\right)}$) to learn features, making it more lightweight and suitable for real-time applications.
3.3.2 Semi-Supervised Pipeline
To facilitate cross-plane image and LiDAR feature synergy training with limited voxel annotations, we adopt the weak-to-strong consistency principle [35] within a semi-supervised learning. The teacher and student models share the same structure. Table 1 lists their data-augmentation strategy.
Table 1: Data augmentation strategy. The number of signs indicates the augmentation strength.
Image LiDAR scan
Brightness HSV space Motion blur Weather sim. Cutout Point dropout Voxelization
Teacher
Student ✓✓✓ ✓✓✓ ✓✓ ✓✓ ✓✓✓ ✓✓✓
We classify image augmentations by their effects on features. Brightness, contrast, and HSV adjustments modify global characteristics and constitute weak augmentation. Weather simulation and motion blur degrade feature clarity and serve as middle augmentation. Cutout removes entire regions of the feature map and thus represents strong augmentation. For LiDAR scans, we vary augmentation strength by adjusting the dropout probability, the dropout number, and the sampling number.
During knowledge distillation, we filter the predictions of teacher model based on confidence scores, which are computed as the maximum channel prediction for each sample. We then concatenate the filtered predictions with the limited voxel annotations to supervise the student model.
3.4 Semantic-Geometric Feature Grid Fusion
In this section, we fuse semantic and geometric feature grids to predict 3D semantic occupancy. Previous works [32, 34] employs attention mechanisms to enable long-range interactions within the feature grid. However, with sequence length ${HWZ}$, 3D DA still incurs substantial computational overhead. Recent work [36] adapts Mamba [17] module for efficient long-range dependency learning in this task. Inspired by this approach and given the spatial alignment of our semantic and geometric feature grids, we propose a Dual Mamba that process semantic and geometric feature grids independently, exchanges subsets of their channels for fusion. Specifically, we lift the feature map ${{\bf{F}}_{I}^{s}}$ to a feature grid via Equ. 2 and combine it with ${{\bf{F}}_{P}^{s}}$ to yield the semantic feature grid ${{\bf{F}}_{f}^{s}}$, Dual mamba then takes ${{\bf{F}}_{f}^{s}}$ and ${{\bf{F}}_{f}^{g}}$ as input and outputs the fused feature grid ${{\bf{F}}_{3d}^{f}}$. The detailed structure of Dual Mamba is illustrated in Appendix A.
GaussTR [11] projects Gaussian feature grids onto the image plane and supervises predictions using pseudo labels from 2D segmentation. Inspired by this approach and aiming to leverage the aligned pseudo label ${{\bf{\bar{s}}}}$ for supervising 3D semantic occupancy prediction, we propose a simple and effective method, termed scatter-accumulated projection, as follows:
${\bf{F}}_{2d}^{f}=\sum\limits_{\bf{p}}{{\bf{F}}_{3d}^{f}}$ (5)
where ${\sum\limits_{\bf{p}}{\left(\cdot\right)}}$ denotes scattered accumulation over index ${{\bf{p}}}$. This operation projects ${{\bf{F}}_{3d}^{f}}$ onto the image plane and aggregates projected feature along each ray. It enables the model to identify the semantic class of each feature grid and to capture occlusion and occupancy states in geometry, thereby facilitating both semantic and geometric supervision. Additionally, this method is agnostic to the specific format of the feature grid, thus offering better scalability.
We use a 3D header and 2D header to predict the 3D semantic occupancy ${{\bf{Y}}}$ and the projected segmentation ${{\bf{y}}\in{{\mathbb{R}}^{h\times w}}}$ from feature grid ${{\bf{F}}_{3d}^{f}}$ and projected feature map ${{\bf{F}}_{2d}^{f}}$.
3.5 Losses and Training Strategy
3.5.1 Losses
During training, we define the loss functions for the semantic branch ${{\ell_{sem}}}$, geometric branch ${{\ell_{geo}}}$, and semantic-geometric feature grid fusion ${{\ell_{fus}}}$ as follows:
$\left\{\begin{array}[]{l}{\ell_{sem}}={\ell_{ce}}\left({{\bf{s}},{\bf{\bar{s}}}}\right)+{\ell_{occ}}\left({{\bf{S}},{\bf{\bar{S}}}}\right)\\ {\ell_{geo}}={\ell_{ce}}\left({{\bf{G}},{\bf{\bar{G}}}}\right)\\ {\ell_{fus}}={\ell_{occ}}\left({{\bf{Y}},{\bf{\bar{Y}}}}\right)+{\ell_{ce}}\left({{\bf{y}},{\bf{\bar{s}}}}\right)\end{array}\right.$ (6)
where ${{\ell_{ce}}}$ denotes the weighted cross-entropy loss, with weights computed from class frequencies. ${{\ell_{occ}}}$ is a comprehensive occupancy loss, following prior work [4, 37, 38]. ${{\bf{\bar{G}}}}$ represents the ground-truth of geometry. We apply ${{\ell_{occ}}\left({{\bf{S}},{\bf{\bar{S}}}}\right)}$ only to voxels containing more than one point, supervising voxel-level semantic learning.
3.5.2 Training Strategy
The training strategy of OccLE is summarized in Algorithm 1. In Phases I and II, we independently train the semantic and geometric branches. In Phase III, all modules are jointly optimized while freezing the geometric branch to prevent its overfitting.
Input: Anotated dataset ${{{\cal D}_{l}}}$, unanotated dataset ${{{\cal D}_{u}}}$, Semantic branch ${{\Theta_{sem}}}$, Geometric branch ${{\Theta_{geo}}}$, Semantic-geometric feature grid fusion module ${{\Theta_{fus}}}$
Output: Trained ${{\Theta_{sem}}}$,${{\Theta_{geo}}}$,${{\Theta_{fus}}}$
Phase I: Train the semantic branch
Train ${{\Theta_{sem}}}$ with dataset ${{{\cal D}_{l}}}$,${{{\cal D}_{u}}}$, loss ${{\ell_{sem}}}$, label ${\{{\bf{\bar{S}}},{\bf{\bar{s}}}\}}$
Phase II: Train the geometric branch
Train the teacher model ${{\Theta_{geo-T}}}$ with dataset ${{{\cal D}_{l}}}$, loss ${{\ell_{sem}}}$, label ${\{{\bf{\bar{G}}}\}}$
Predict ${\{{{\bf{\bar{G}}}_{T}}\}}$ for dataset ${{{\cal D}_{u}}}$ via ${{\Theta_{geo-T}}}$
Train the student model ${{\Theta_{geo-S}}}$ with dataset ${{{\cal D}_{l}}}$,${{{\cal D}_{u}}}$, loss ${{\ell_{sem}}}$, label ${\{{\bf{\bar{G}}}\}}$,${\{{{\bf{\bar{G}}}_{T}}\}}$
Phase III: Train the semantic-geometric feature grid fusion module
Freeze ${{\Theta_{geo-S}}}$; Train ${{\Theta_{fus}}}$ and ${{\Theta_{sem}}}$ with dataset ${{{\cal D}_{l}}}$,${{{\cal D}_{u}}}$, loss ${{\ell_{fus}}}$,${{\ell_{sem}}}$, label ${\{{\bf{\bar{Y}}},{\bf{\bar{s}}}\}}$
Algorithm 1 The training strategy of OccLE
4 Experiment
4.1 Dataset and Metric
4.1.1 Dataset
We evaluate OccLE on the SemanticKITTI [27], which includes 22 outdoor driving scenarios: 10 sequences for training, 1 for validation, and 11 for testing. The dataset annotates scenes of size ${51.2m\times 51.2m\times 6.4m}$ with a voxel size of ${0.2m\times 0.2m\times 0.2m}$ for 20 semantic classes (19 semantics + 1 free). To simulate a label-efficient learning, we use only 10% of the voxel annotations from the training set, obtained via interval sampling.
4.1.2 Metric
There are two domain metrics: intersection over union (IoU) and mean intersection over union (mIoU) across 19 classes. IoU measures the overall scene completion quality, while mIoU evaluates the quality of semantic segmentation for each class. We use IoU as the primary metric for the geometric branch and mIoU as the primary metric for the semantic branch and semantic-geometric feature grid fusion.
4.2 Implementation Details
[…]
4.3 Main Result
We present a quantitative comparison on the SemanticKITTI validation set in Table 2. OccLE obtains competitive performance, achieving 16.59 % mIoU and 39.96 % IoU. It surpasses all fully supervised camera-based methods (HASSC [40] with 40.66 % mIoU) and approaches the performance of fully supervised multi-modal methods (OccLoff [22] with 22.62 % mIoU). As illustrated in Figure 4, OccLE achieves strong performance in complex scenes, accurately reconstructing buildings and vegetation, and distinguishing small objects such as traffic signs and poles.
Table 2: Quantitative results on SemanticKITTI validation set. Bold and underline represent the best and second best results, respectively.
Method Inp. Sup. IoU mIoU
car (3.92%)
bicycle (0.03%)
motorcycle (0.03%)
truck (0.16%)
other-veh. (0.20%)
person (0.07%)
bicyclist (0.07%)
motorcyclist (0.05%)
road (15.30%)
parking (1.12%)
sidewalk (11.13%)
other-grnd.(0.56%)
building (14.10%)
fence (3.90%)
vegetation (39.3%)
trunk (0.51%)
terrain (9.17%)
pole (0.29%)
traf.-sign (0.08%)
TPVFormer [19] C Full 35.61 11.36 23.81 0.36 0.05 8.08 4.35 0.51 0.89 0.00 56.50 20.60 25.87 0.85 13.88 5.94 16.92 2.26 30.38 3.14 1.52
OccFormer [14] C Full 36.50 13.46 25.09 0.81 1.19 25.53 8.52 2.78 2.82 0.00 58.85 19.61 26.88 0.31 14.40 5.61 19.63 3.93 32.62 4.26 2.86
VoxFormer [4] C Full 44.15 13.35 25.64 1.28 0.56 7.26 7.81 1.93 1.97 0.00 53.57 19.69 26.52 0.42 19.54 7.31 26.10 6.10 33.06 9.15 4.94
PanoSSC [41] C Full 34.94 11.22 19.63 0.63 0.36 14.79 6.22 0.87 0.00 0.00 56.36 17.76 26.40 0.88 14.26 5.72 16.69 1.83 28.05 1.94 0.70
H2GFormer [32] C Full 44.69 14.29 28.21 0.95 0.91 6.80 9.32 1.15 0.10 0.00 57.00 21.74 29.37 0.34 20.51 7.98 27.44 7.80 36.26 9.88 5.81
OctreeOcc [42] C Full 44.71 13.12 28.07 0.64 0.71 16.43 6.03 2.25 2.57 0.00 55.13 18.68 26.74 0.65 18.69 4.01 25.26 4.89 32.47 3.72 2.36
SGN [38] C Full 46.21 15.32 33.31 0.61 0.46 6.03 9.84 0.47 0.10 0.00 59.10 19.05 29.41 0.33 25.17 9.96 28.93 9.58 38.12 13.25 7.32
Symphonies [37] C Full 41.44 13.44 27.23 1.44 2.28 15.99 9.52 3.19 8.09 0.00 55.78 14.57 26.77 0.19 18.76 6.18 24.50 4.32 28.49 8.99 5.39
HASSC [40] C Full 44.55 15.88 30.64 1.20 0.91 23.72 7.77 1.79 2.47 0.00 62.75 20.20 32.40 0.51 22.90 8.67 26.47 7.14 38.10 9.00 5.23
RenderOcc [6] C Full - 12.87 24.90 0.37 0.28 6.03 3.66 1.91 3.11 0.00 57.2 16.11 28.44 0.91 18.18 9.10 26.23 4.87 33.61 6.24 3.38
OccLoff [22] C+L Full - 22.62 46.44 2.08 3.91 20.38 8.72 3.88 4.35 0.00 66.25 21.07 43.51 0.57 41.23 15.86 41.20 20.06 46.21 27.60 16.47
OccFusion [5] C+L Full 58.68 21.92 45.62 2.96 3.51 20.05 8.76 3.16 4.37 0.00 65.67 23.08 36.33 0.00 39.09 15.70 45.53 19.37 45.53 27.57 15.21
OccGen [23] C+L Full 36.87 13.74 26.83 1.60 2.53 15.49 12.83 3.20 3.37 0.00 61.28 20.42 28.30 0.43 14.49 6.94 20.04 3.94 32.44 4.11 2.77
RenderOcc [6] C Self - 8.24 14.83 0.42 0.17 2.47 1.78 0.94 3.20 0.00 43.64 12.54 19.10 0.00 11.59 4.71 17.61 1.48 20.01 1.17 0.88
OccLE(Ours) C+L 10% 40.60 16.59 37.27 2.76 4.39 0.90 10.88 2.52 0.09 0.00 55.79 19.96 28.85 0.14 24.56 12.53 35.19 16.98 35.29 17.16 10.01
Figure 4: Qualitative results on the SemanticKITTI validation set. OccF., VoxF., SGN, Sym., and HASSC represent the prediction results from [14], [4], [38], [37], and [40], respectively. GT denotes the ground truth.
4.4 Ablation Studies
We conduct all ablation studies on the validation set of the SemanticKITTI dataset. We focus on three aspects: semantic branch, geometric branch, and overall pipeline.
4.4.1 The Semantic Branch
To evaluate the extractor and pseudo label in the semantic branch, for comparison, we adopt ESPNetv2 [43] and MobileNet [44] as ${{{\cal G}_{s2d}}}$, employ 1-layer and 2-layer versions of sparse UNet3D as ${{{\cal G}_{s3d}}}$, and train using unaligned pseudo labels. As depicted in Table 3, 2D extractors designed for edge device suffer a notable performance degradation, whereas our method achieves the highest mIoU(${{\bf{s}}}$) while retaining remaining lightweight. Increasing the depth of the sparse UNet3D architecture improves mIoU(${{\bf{S}}}$) without substantially increasing computational cost. Moreover, aligned pseudo labels yield a significant mIoU gain by providing supervision for unaligned classes.
Table 3: Ablation study of the semantic branch. The unit of Inf. Time is in milliseconds.
${{{\cal G}_{s2d}}}$ Inf. Time mIoU(${{\bf{s}}}$) ${{{\cal G}_{s3d}}}$ Inf. Time mIoU(${{\bf{S}}}$)
ESPNetv2 4.38 28.71 1-layer 41.77 12.38
MobileNet 6.10 48.76 2-layer 43.22 13.30
Ours (w/o align.) 6.87 41.89 Ours (w/o align.) 44.59 9.89
Ours (w/ align.) 6.87 51.51 Ours (w/ align.) 44.59 13.94
4.4.2 The Geometric Branch
To evaluate the effects of model architecture and training strategy in the geometric branch, we design a baseline that employs Conv3D and 3D DA for feature extraction and interaction, and compare our semi-supervision against full supervision. As depicted in Table 4(a), our cross-plane image and LiDAR feature synergy only has a minor precision drop while significantly reducing module inference time from 138.62 ms to 26.08 ms. Our data augmentation strategy boosts student model performance as augmentation degree increases. Under semi-supervision, our student model reaches 53.60 IoU, closely approaching the fully supervised performance of 56.77 IoU.
Table 4: Ablation studies for (a) geometric branch and (b) overall pipeline. “weak,” “medium,” and “strong” denote the degrees of data augmentation, -S and -T represents the student and teacher models. ‘Sem.’, ‘Geo.’, and ‘Proj.’ denote the semantic branch, geometric branch, and supervising with scatter-accumulated projection, respectively.
(a) Geometric branch.
Sup. Weak Middle Strong Model IoU
Full Ours-T 56.77
10% Baseline 52.79
10% Ours-T 52.66
10% Ours-S 52.86
10% Ours-S 52.88
10% Ours-S 53.60
(b) Overall pipeline.
Sup. Sem. Geo. Proj. Model mIoU IoU
Full Ours 20.32 51.73
10% Baseline 9.64 22.30
10% Ours 11.84 27.46
10% Ours 8.08 17.52
10% Ours 11.55 18.52
10% Ours 16.59 40.60
4.4.3 Overall Pipeline
As depicted in Table 4(b), compared to the baseline, our method increases mIoU from 9.64 to 16.59 using only 10% voxel annotations, closely approaching the fully supervised performance of 20.32 mIoU. Furthermore, ablating any module in OccLE results in a precision drop, demonstrating the essential role of these three modules in label-efficient learning. Additional ablation studies are presented in the Appendix C.2.
5 Conclusion
In this paper, we propose OccLE, a label-efficient 3D semantic occupancy prediction that leverages limited voxel annotations while preserving high performance. OccLE incorporates three key components: (1) distilling 2D foundation model to predict aligned pseudo label for supervising 2D and 3D semantic learning, (2) cross-plane image and LiDAR feature synergy for efficient geometry learning under semi-supervision, and (3) semantic-geometric feature grid fusion with Dual Mamba and scatter-accumulated projection for supervising unannotated regions. Experimental results demonstrate that OccLE achieves an mIoU of 16.59% on the SemanticKITTI validation set using only 10% of the voxel annotations.
Appendix A Dual Mamba Structure
The detailed structure of Dual Mamba is illustrated in Fig. A1. The model employs a U-Net backbone with four downsampling stages and four upsampling stages. Each stage incorporates a dual Mamba block. Within a dual Mamba block, two parallel branches compute positional embeddings for the two inputs and then exchange their feature channels prior to reordering. We use a Hilbert curve [45] to linearize the 3D representation into a 1D sequence, apply a Mamba block to capture long-range dependencies, and finally restore the three-dimensional structure.
Figure A1: The detailed structure of Dual Mamba. It comprises four stages and employs two parallel branches to process the inputs at each stage.
Appendix B Aligned Label Generation
We employ dataset-specific and open-vocabulary segmentation 2D foundation models to generate aligned pseudo labels. The class mappings are listed in Table B.1. We utilize SAM2 [8] to produce trunk class labels and integrate them into the MSeg [28] outputs. This example illustrates our aligned pseudo label generation process, which can scale to any dataset with different class taxonomies.
Table B.1: The aligned label generation.
SemantiKITTI Image Segmentation SemantiKITTI Image Segmentation
Label Class Label Class Label Class Label Class
0 free 12 other-grnd. 31 road_barrier
1 car 176 car 32 mailbox
2 bicycle 175 bicycle 137 fire_hydrant
3 motorcycle 178 motorcycle 191 wall
13 building 35 building
4 truck 182 truck 14 fence 144 fence
5 other.-veh. 177 autorickshaw 15 vegetation 131 road_barrier
180 bus 174 vegetation
181 train 16 trunk SAM2 trunk
183 trailer 17 terrain 102 terrain
185 slow_wheeled_object 18 pole 143 pole
6 person 125 person 130 streetlight
7 bicyclist 126 rider_other 145 railing_banister
127 bicyclist 146 guard_rail
8 motorcyclist 128 motorcyclist 162 column
9 road 98 road 19 traf.-sign. 135 traffic_sign
10 parking 138 parking_meter 136 traffic_light
11 sidewalk 100 sidewalk_pavement
Appendix C Supplementary Experimental Results
C.1 Supplementary Ablation Study
We evaluate some experimental setting in Table C.2. To assess the impact of semi-supervision in the geometric branch on the overall pipeline, we incorporated a teacher model of the geometric branch into final joint training. The results indicate that mIoU and IoU decrease by 0.42 and 1.74, respectively.
We compare our scatter-accumulated projection method with an alternative approach that weights features by the distance between each voxel and the camera. The latter method yields a 3.21 mIoU drop. This result demonstrates that the scatter-accumulated projection is simple yet effective in supervising all voxels with aligned pseudo labels, whereas the distance-weighted projection reduces supervision quality for distant voxels.
We assess the necessity of using limited voxel annotations rather than relying solely on aligned pseudo labels. Our experiments indicate that relying solely on aligned pseudo labels results in a marked decline in precision, signaling pattern collapse during training when voxel annotations are absent. These results confirm that voxel annotations are essential for guiding the model toward an effective fusion of semantic and geometric features.
Table C.2: Supplementary ablation study.
Model mIoU IoU
w/ geometric teacher model 16.17 38.86
w/ weighted projection 13.38 40.40
w/o limited voxel annotation 0.37 7.84
Ours 16.59 40.60
C.2 Supplementary Qualitative Comparison
We show additional qualitative results of OccLE on the SemanticKITTI validation dataset in Figure C1. OccLE maintains strong performance across varied scenes, delivering effective scene completion and precise class classification.
Figure C1: The qualitative results of OccLE on SemanticKITTI validation dataset.
Appendix D Limitation
In this study, we simulate label-efficient learning for 3D semantic occupancy prediction by uniformly sampling 10% of voxel annotations across all scenes in the SemanticKITTI dataset. This strategy ensures the generalization across diverse scenes. However, concentrating limited voxel annotations within a few scenes may lead to overfitting, thereby diminishing the robustness of OccLE in unfamiliar scenes. Consequently, for real-world autonomous driving applications, it is advisable to distribute limited voxel annotations across a wide range of scenes, enhancing its performance in varied and unseen environments.

### v2

@@
 We fuse semantic-geometric feature grids through Dual Mamba and incorporate a scatter-accumulated projection to supervise unannotated prediction with aligned pseudo labels.
-Experiments show that OccLE achieves competitive performance with only 10% of voxel annotations, reaching a mIoU of 16.59% on the SemanticKITTI validation set.
-Figure 1: Label-efficient 3D semantic occupancy prediction aims to achieve high performance using limited voxel annotations and aligned pseudo label.
-We propose OccLE, a novel learning paradigm that decouples semantic and geometric learning and fuse their feature grids for the final prediction.
-Keywords Autonomous Driving  $\cdot$ 3D Semantic Occupancy Prediction  $\cdot$ Label-Efficient Learning
+Experiments show that OccLE achieves competitive performance with only 10% of voxel annotations on the SemanticKITTI and Occ3D-nuScenes datasets.
+Keywords— Autonomous Driving 3D Semantic Occupancy Prediction Label-Efficient Learning
 1 Introduction
 3D perception task is foundational cornerstone for autonomous driving systems.
-Among various perception methods, 3D semantic occupancy prediction [1, 2, 3] has garnered significant attention for providing intuitive and efficient scene understanding for downstream tasks.
-This task aims to estimate the occupancy status and semantic label of each voxel in a 3D grid, given input from 2D images, 3D LiDAR, or a combination of both.
+Among various perception methods, 3D semantic occupancy prediction [36, 24] has garnered significant attention for providing intuitive and efficient scene understanding for downstream tasks.
+This task estimates the occupancy status and semantic label of each voxel in a 3D grid, given input from 2D images, LiDAR, or a combination of both.
 Previous studies [22, 30] proposed numerous supervised learning paradigms with diverse modal inputs.
@@
 Even with pre-annotated 3D labels generated using auto-labeling assistants, one hour of human labor just only annotates 10 frames [31, 14].
-This issue limits the scalability of supervised methods for real-world deployment, especially in terms of robustness.
+This issue limits the scalability and robustness of supervised methods in real-world deployment.
+Figure 1: Label-efficient 3D semantic occupancy prediction aims to achieve high performance using limited voxel annotations and aligned pseudo label.
+We propose OccLE, a novel learning paradigm that decouples semantic and geometric learning and fuse their feature grids for the final prediction.
 Recent advancements in 3D semantic occupancy prediction have shifted towards a label-free paradigm.
 These self-supervised methods leverage existing vision foundation models to generate various pseudo labels for auxiliary supervision, such as image semantic segmentation [32], depth information [48], and LiDAR segmentation [57].
-To bridge the 3D representation with 2D pseudo label, specific 3D representation formats have been proposed to facilitate volume rendering, including signed distance functions[7] and 3D Gaussians [11].
+To bridge the 3D representation with 2D pseudo label, they proposed specific 3D representation formats to facilitate volume rendering, including signed distance functions [14] and 3D Gaussians [18].
 The core strategy of these self-supervised methods is to supplement pseudo semantic supervision.
 They project LiDAR segmentation annotations from the same dataset into 2D pseudo labels [3, 31].
-However, this pseudo labels generation still is expensive and provides only sparse supervision.
+However, this pseudo labels still is expensive and provides only sparse supervision.
 They also employ a 2D vision foundation model to generate 2D pseudo labels [14], which suffers from class misalignment.
@@
 1, which uses limited voxel annotations while maintaining high performance.
-We design the label-efficient learning paradigm: taking image and LiDAR scan inputs, fully decoupling semantic and geometric feature learning, fusing their feature grids for 3D semantic occupancy prediction, and supervising with limited voxel annotations and aligned pseudo labels.
+We design the label-efficient learning paradigm: taking image and LiDAR inputs, fully decoupling semantic and geometric feature learning, fusing their feature grids for 3D semantic occupancy prediction, and supervising with limited voxel annotations and aligned pseudo labels.
 Specifically, OccLE consists of the semantic branch, the geometric branch, and semantic-geometric feature grid fusion.
@@
 • We introduce a label-efficient 3D semantic occupancy prediction task and a learning paradigm that decouples semantic and geometric learning, followed by their synergistic integration.
-The model is supervised using limited voxel annotations and aligned pseudo labels to achieve high performance.
-• We present three techniques: distilling 2D foundation model for 2D and 3D semantic learning; semi-supervised geometry learning using cross-plane image and LiDAR feature synergy; and semantic-geometric feature grids fusion with auxiliary supervision by scatter-accumulated projection.
-• Experimental results show that OccLE achieves competitive performance compared to fully supervised methods, attaining 16.59 % mIoU on the SemanticKITTI validation set using only 10 % of the voxel annotations.
+The model is supervised with limited voxel annotations and aligned pseudo labels to achieve high performance.
+• We present three techniques: distilling 2D foundation model for 2D and 3D semantic learning; semi-supervised geometry learning using cross-plane image and LiDAR feature synergy; and semantic-geometric feature grids fusion with scatter-accumulated projection auxiliary supervision.
+• Experimental results show that OccLE achieves competitive performance compared to fully supervised methods, attaining 16.59 % and 27.53% mIoU on the SemanticKITTI validation set and Occ3D-nuScenes using only 10 % of the voxel annotations.
 2 Related Works
@@
 For camera-based 3D semantic occupancy prediction, a key challenge lies in lifting 2D to 3D representations.
-Previous works [13, 14, 15] utilize calibration and depth information to establish explicit backwards projections.
-ViewFormer [16] proposes a learning-first attention mechanism to aggregate multi-view features.
-Some studies [17, 18] represent scenes using sparse 3D semantic Gaussians.
+Previous works [53, 42] utilize calibration and depth information to establish explicit backwards projections.
+ViewFormer [21] proposes a learning attention mechanism to aggregate multi-view features.
+Some studies [8, 16] represent scenes using 3D Gaussians.
 TPVFormer [13] introduces three perpendicular views (TPV) to enhance the interaction of image features.
-Since dimensional transformation based on depth estimation is ill-posed for camera-based methods, multi-modal method incorporates LiDAR and radar inputs to obtain accurate distance measurements.
+Since dimensional transformation based on estimated depth is ill-posed for camera-based methods, multi-modal method incorporates LiDAR and radar inputs to obtain accurate distance measurements.
 Its essential research is modal fusion.
@@
 MetaOcc [46] learn global and local alignment to enhance fused representations.
-Some studies [21, 22] advance fusion by employing point-to-point queries from 3D to 2D.
+Some studies [50, 52] employ point-to-point queries from 3D to 2D.
 OccGen [38] employs a diffusion model to refine the fused features.
-Building on prior work, OccLE takes camera images and a LiDAR scans as inputs, applies backward projection to derive 3D feature grids from 2D image feature maps, and proposes a lightweight yet effective fusion strategy in the geometry learning termed cross-plane image and LiDAR feature synergy.
+Building on prior work, OccLE takes camera images and LiDAR scans as inputs, applies backward projection to derive 3D feature grids from 2D image feature maps, and proposes a lightweight fusion strategy for geometry learning, termed cross-plane image and LiDAR feature synergy.
 2.2 Label-efficient learning
@@
 Semi-supervised learning [2] combines a small labeled dataset with abundant unlabeled data to improve model generalization, while self-supervised learning [9] generates supervision directly from unlabeled data through pretext tasks.
-In label-efficient learning for 3D semantic occupancy prediction, self-supervision methods render 3D feature into 2D representations, and supervise with labels from other tasks.
+In label-efficient learning for 3D semantic occupancy prediction, self-supervision methods render 3D feature into 2D format, and supervise with labels from other tasks.
 To facilitate rendering, SelfOcc [14] introduces a signed distance function-based representation, while GaussTR [18] proposes a 3D Gaussian-based representation.
-For auxiliary supervision, OccFlowNet [12] utilizes projected semantic point clouds, whereas RenderOcc [6] relies on depth estimation and image segmentation.
+For auxiliary supervision, OccFlowNet [3] utilizes projected LiDAR semantic, whereas RenderOcc [31] relies on estimated depth and image segmentation.
 Inspired by the above, OccLE utilizes the pseudo label from image segmentation as a key supervision to enable label-efficient learning.
@@
 Finally, we fuse semantic and geometric feature grids via Dual Mamba and supervise the unanotated prediction with aligned pseudo label using scatter-accumulated projection.
-We aim to take multi-frame camera images ${\left\{{{{\bf{I}}_{t-i}}}\right\}_{i=0}^{N-1}\in{{\mathbb{R}}^{h\times w}}}$, multi-frame point clouds ${\left\{{{{\bf{P}}_{t-i}}}\right\}_{i=0}^{N-1}\in{{\mathbb{R}}^{n\times 4}}}$, camera calibration ${\left\{{{\bf{K}}{\rm{,}}{\bf{T}}}\right\}}$ as inputs to predict the 3D semantic occupancy ${{{\bf{Y}}_{t}}\in{\left\{{{c_{0}},{c_{1}},\cdots,{c_{M-1}}}\right\}^{H\times W\times Z}}}$ at timestep ${t}$.
-To enable label-efficient training, we supervise the model using only limited voxel annotations (e.g., 10% ${\left\{{{{{\bf{\bar{Y}}}}_{t}}}\right\}}$ of all samples).
+We aim to take multi-frame camera images $\left\{{{{\bf{I}}_{t-i}}}\right\}_{i=0}^{N-1}\in{{\mathbb{R}}^{h\times w\times 3}}$}, multi-frame point clouds $\left\{{{{\bf{P}}_{t-i}}}\right\}_{i=0}^{N-1}\in{{\mathbb{R}}^{n\times 4}}$}, camera calibration $\left\{{{\bf{K}}{\rm{,}}{\bf{T}}}\right\}$} as inputs to predict the 3D semantic occupancy ${{\bf{Y}}_{t}}\in{\left\{{{c_{0}},{c_{1}},\cdots,{c_{M-1}}}\right\}^{H\times W\times Z}}$} at timestep t{t}.
+To enable label-efficient training, we supervise the model training with limited voxel annotations (e.g., 10% $\left\{{{{{\bf{\bar{Y}}}}_{t}}}\right\}$} of all samples).
 Here, ${\bf{K}}{\rm{,}}{\bf{T}}$} represent the intrinsic and extrinsic camera matrices, respectively; $\left\{{{{{\bf{\bar{Y}}}}_{t}}}\right\}$} is the ground truth of 3D semantic occupancy prediction.
-${N}$ denotes frame number; ${M}$ denotes class number, ${h}$ and ${w}$ denote the height and width of camera image; ${H}$, ${W}$, and ${Z}$ represent the length, width, height of voxel grid.
+N{N} denotes frame number; M{M} denotes class number, h{h} and w{w} denote the height and width of camera image; H{H}, W{W}, and Z{Z} represent the length, width, height of voxel grid.
 Our method is formulated as ${{\bf{Y}}_{t}}=\Theta\left({\left\{{{{\bf{I}}_{t-i}}}\right\}_{i=0}^{N-1},\left\{{{{\bf{P}}_{t-i}}}\right\}_{i=0}^{N-1},{\bf{K}}{\rm{,}}{\bf{T}}}\right)$}, where $\Theta$} is our proposed OccLE.
-For clarity, we omit the timestep subscript ${t}$ for all variables in the following description.
+For clarity, we omit the timestep subscript t{t} for all variables in the following description.
 The overview of OccLE is illustrated in Fig.
@@
 In the semantic-geometric feature grid fusion module, we employ Dual Mamba to fuse feature grids and employ scatter-accumulated projection to supervise unanotated predictions with aligned pseudo labels.
-The semantic branch, the geometric branch, and the semantic-geometric feature grid fusion will be discussed in Sec.
-3.2, Sec.
-3.3, and Sec.
-3.4, respectively.
 3.2 Distill 2D Foundation Models for 2D and 3D Semantic Learning
@@
 To mitigate this requirement, we notice that the image semantic segmentation maps the feature map to pixel semantic label and benefits from large annotated datasets.
-Therefore, we aim to distill 2D foundation model for 2D and 3D semantic learning, where the 2D foundation model predicts aligned pseudo labels to supervise the semantic learning of image and LiDAR scan inputs.
+Therefore, we aim to distill 2D foundation model for 2D and 3D semantic learning, where the 2D foundation model predicts aligned pseudo labels to supervise the semantic learning of image and LiDAR inputs.
 2D foundation models may not share the same class definitions as 3D semantic occupancy prediction.
@@
 We adapt MSeg universal class classification to the task-specific classes, input the unaligned class name prompts into SAM2, and merge outputs using Boolean operations to generate aligned pseudo labels $\bf{\bar{s}}$}}.
-We employ ${{\bf{\bar{s}}}}$ to fully supervise semantic learning of images and LiDAR scans, enabling lightweight deployment via distillation.
+We employ $\bf{\bar{s}}$}} to fully supervise semantic learning of images and LiDAR, enabling lightweight deployment via distillation.
 Specifically, we employ a UNet [34] ${\cal G}_{s2d}$}} to extract features map ${\bf{F}}_{I}^{s}$} from the image $\bf{I}$}}.
-We adopt the voxelization in VoxelNet [30] and utilize a sparse UNet3D [31] ${{{\cal G}_{s3d}}}$ to extract features grid ${{\bf{F}}_{P}^{s}}$ from the LiDAR scan ${{\bf{P}}}$.
+We adopt the voxelization in VoxelNet [56] and utilize a sparse UNet3D [44] ${\cal G}_{s3d}$}} to extract features grid ${\bf{F}}_{P}^{s}$} from the LiDAR $\bf{P}$}}.
 The mathematical expression is given by:
-[…]
+$\left\{\begin{aligned} \mathbf{F}_{I}^{s}&=\mathcal{G}_{s2}(\mathbf{I})\\ \mathbf{F}_{P}^{s}&=\mathcal{G}_{s3}(\mathcal{V}(\mathbf{P}))\end{aligned}\right$. (1)
 where $\cal V$}} refers to the voxelization process, which randomly samples up to 35 points per voxel. Each point includes its position, intensity, and the offset from the voxel center.
@@
 ${{\bf{\bar{S}}}}={\cal F}\left({{\bf{\bar{s}}},{\bf{p}}}\right),{\bf{p}}={\bf{K}}\left({{\bf{T}}\cdot{\bf{P}}}\right)$ (2)
-where ${{\cal F}}$ denotes the sampling function, ${{\bf{P}}}$ denote the voxel grid coordinates, ${{\bf{p}}}$ represent the corresponding projected coordinates on the image plane.
+[…]
 3.3 Semi-Supervised Geometry Learning
@@
 Because depth estimation is ill-posed, reducing voxel annotations exacerbates uncertainty in this sequential prediction.
-Since depth is sourced from the LiDAR scan projected on the image plane, we directly use image and LiDAR scan inputs, propose a lightweight and modality-complementary geometry learning module, termed cross-plane image and LiDAR feature synergy, and employ a semi-supervision to learn an accurate geometric feature grid with limited voxel annotations.
+Since depth is sourced from the LiDAR projected on the image plane, we directly use image and LiDAR inputs, propose a lightweight and modality-complementary geometry learning module, termed cross-plane image and LiDAR feature synergy, and employ a semi-supervision to learn an accurate geometric feature grid with limited voxel annotations.
 3.3.1 Cross-Plane Image and LiDAR Feature Synergy
-First, we examine the characteristics of the image and LiDAR scan inputs.
+First, we examine the characteristics of the image and LiDAR inputs.
 As shown by the red boxes in Fig.
 3a and Fig.
-3b, the frontal-view LiDAR scan is sparse and omits information in the upper region, whereas the image offers continuous context.
-In the bird’s-eye view (BEV), the image cannot resolve distances along each ray, while the LiDAR scan provides accurate range measurements.
-In summary, the image is reliable on the ${yz}$-plane and the LiDAR scan is reliable along the ${x}$-axis in world coordinates.
-Motivated by their inherency and inspired by [33], we project each their feature grid into its optimal plane in the TPV and fuse them via multiplicative integration.
+3b, the frontal-view LiDAR is sparse and omits information in the upper region, whereas the image offers continuous context.
+In the bird’s-eye view (BEV), the image cannot resolve distances along each ray, while the LiDAR provides accurate range measurements.
+In summary, the image is reliable on the yz{yz}-plane and the LiDAR is reliable along the x{x}-axis in world coordinates.
+Motivated by their inherency and inspired by [51], we propose using TPV to achieve a synergistic and lightweight multi-modal features fusion by projecting features onto their optimal TPV planes and applying multiplicative integration.
 Figure 3: Illustration of geometry learning. (a) Frontal view feature comparison. (b) BEV view feature comparison. (c) The cross-plane image and LiDAR feature synergy.
@@
 3c.
-We voxelize multi-frame LiDAR scans using the VoxelNet [30] and process them with a sparse 3D encoder to obtain LiDAR feature grids.
+We voxelize multi-frame LiDAR using the VoxelNet [56] and process them with a sparse 3D encoder to obtain LiDAR feature grids.
 In parallel, we extract feature maps from multi-frame images using a 2D encoder and lift into the feature grid defined in Equ.
@@
 We fuse the multi-frame feature grids into a single grid via a Conv1D layer.
-For the LiDAR feature grid, we stack along the ${y}$ and ${z}$ axes and apply a Conv2D layer to produce feature maps ${{\bf{F}}_{I,xz}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{Z_{1}}}}}$ and ${{\bf{F}}_{I,xy}^{g}\in{{\mathbb{R}}^{{W_{1}}\times{Z_{1}}}}}$.
-For the image feature grid, we stack along the ${x}$-axis and apply a Conv2D layer to generate feature map ${{\bf{F}}_{P,yz}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}}$.
-To enhance ${{\bf{F}}_{P,xy}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}}$, we apply self-attention (SA) along the ${y}$ and ${z}$ axes, yielding feature maps ${{\bf{F}}_{P,yzy}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}}$, ${{\bf{F}}_{P,yzz}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}}$ as follows:
-$\left\{\begin{array}[]{l}{\bf{F}}_{P,yzy}^{g}={\bf{F}}_{P,yz}^{g}+SA\left({{\bf{F}}_{P,yz}^{g}}\right)\\ {\bf{F}}_{P,yzz}^{g}={\bf{F}}_{P,yz}^{g}+SA{\left({{\bf{F}}{{{}_{P,yz}^{g}}^{T}}}\right)^{T}}\end{array}\right.$ (3)
+For the LiDAR feature grid, we stack along the y{y} and z{z} axes and apply a Conv2D layer to produce feature maps ${\bf{F}}_{I,xz}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{Z_{1}}}}$} and ${\bf{F}}_{I,xy}^{g}\in{{\mathbb{R}}^{{W_{1}}\times{Z_{1}}}}$}.
+For the image feature grid, we stack along the x{x}-axis and apply a Conv2D layer to generate feature map ${\bf{F}}_{P,yz}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}$}.
+To enhance ${\bf{F}}_{P,xy}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}$}, we apply self-attention (SA) along the y{y} and z{z} axes, yielding feature maps ${\bf{F}}_{P,yzy}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}$}, ${\bf{F}}_{P,yzz}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}$} as follows:
+$\left\{\begin{array}[]{l}{\bf{F}}_{P,yzy}^{g}={\bf{F}}_{P,yz}^{g}+SA\left({{\bf{F}}_{P,yz}^{g}}\right)\\ {\bf{F}}_{P,yzz}^{g}={\bf{F}}_{P,yz}^{g}+SA{\left({{\bf{F}}{{{}_{P,yz}^{g}}^{T}}}\right)^{T}}\end{array}\right$.
+(3)
 We fuse the image and LiDAR feature maps via matrix multiplication in the TPV.
@@
 $\left\{\begin{array}[]{l}{\bf{F}}_{{f_{1}}}^{g}={\bf{F}}_{I,xz}^{g}\times{\bf{F}}_{P,yzz}^{g}\\ {\bf{F}}_{{f_{2}}}^{g}={\bf{F}}_{I,xy}^{g}\times{\bf{F}}_{P,yzy}^{g}\end{array}\right$. (4)
-We add ${{\bf{F}}_{{f_{1}}}^{g}}$ and ${{\bf{F}}_{{f_{2}}}^{g}}$ to generate the geometric feature grid ${{\bf{F}}_{{f}}^{g}}$ and predict geometry ${{\bf{G}}\in{{\mathbb{R}}^{H\times W\times Z}}}$ via a 3D header.
+We add ${\bf{F}}_{{f_{1}}}^{g}$} and ${\bf{F}}_{{f_{2}}}^{g}$} to obtain the geometric feature grid ${\bf{F}}_{{f}}^{g}$} and predict geometry ${\bf{G}}\in{{\mathbb{R}}^{H\times W\times Z}}$} via a 3D header.
 In the proposed workflow, feature grids are converted into feature maps by stacking and subsequently reconstructed into feature grids during the final fusion.
@@
 Table 1: Data augmentation strategy. The number of signs indicates the augmentation strength.
-Image LiDAR scan
+Image LiDAR
 Brightness HSV space Motion blur Weather sim. Cutout Point dropout Voxelization
@@
 Cutout removes entire regions of the feature map and thus represents strong augmentation.
-For LiDAR scans, we vary augmentation strength by adjusting the dropout probability, the dropout number, and the sampling number.
+For LiDAR, we vary augmentation strength by adjusting the dropout probability, the dropout number, and the sampling number.
 During knowledge distillation, we filter the predictions of teacher model based on confidence scores, which are computed as the maximum channel prediction for each sample.
@@
 Previous works [41, 43] employs attention mechanisms to enable long-range interactions within the feature grid.
-However, with sequence length ${HWZ}$, 3D DA still incurs substantial computational overhead.
+However, with sequence length HWZ{HWZ}, 3D DA still incurs substantial computational overhead.
 Recent work [20] adapts Mamba [8] module for efficient long-range dependency learning in this task.
@@
 Additionally, this method is agnostic to the specific format of the feature grid, thus offering better scalability.
-We use a 3D header and 2D header to predict the 3D semantic occupancy ${{\bf{Y}}}$ and the projected segmentation ${{\bf{y}}\in{{\mathbb{R}}^{h\times w}}}$ from feature grid ${{\bf{F}}_{3d}^{f}}$ and projected feature map ${{\bf{F}}_{2d}^{f}}$.
-3.5 Losses and Training Strategy
-3.5.1 Losses
+We use a 3D and 2D header to predict the 3D semantic occupancy $\bf{Y}$}} and the projected segmentation ${\bf{y}}\in{{\mathbb{R}}^{h\times w}}$} from feature grid ${\bf{F}}_{3d}^{f}$} and projected feature map ${\bf{F}}_{2d}^{f}$}.
+3.5 Losses
 During training, we define the loss functions for the semantic branch $\ell_{sem}$}}, geometric branch $\ell_{geo}$}}, and semantic-geometric feature grid fusion $\ell_{fus}$}} as follows:
-$\left\{\begin{array}[]{l}{\ell_{sem}}={\ell_{ce}}\left({{\bf{s}},{\bf{\bar{s}}}}\right)+{\ell_{occ}}\left({{\bf{S}},{\bf{\bar{S}}}}\right)\\ {\ell_{geo}}={\ell_{ce}}\left({{\bf{G}},{\bf{\bar{G}}}}\right)\\ {\ell_{fus}}={\ell_{occ}}\left({{\bf{Y}},{\bf{\bar{Y}}}}\right)+{\ell_{ce}}\left({{\bf{y}},{\bf{\bar{s}}}}\right)\end{array}\right.$ (6)
+$\left\{\begin{array}[]{l}{\ell_{sem}}={\ell_{ce}}\left({{\bf{s}},{\bf{\bar{s}}}}\right)+{\ell_{occ}}\left({{\bf{S}},{\bf{\bar{S}}}}\right)\\ {\ell_{geo}}={\ell_{ce}}\left({{\bf{G}},{\bf{\bar{G}}}}\right)\\ {\ell_{fus}}={\ell_{occ}}\left({{\bf{Y}},{\bf{\bar{Y}}}}\right)+{\ell_{ce}}\left({{\bf{y}},{\bf{\bar{s}}}}\right)\end{array}\right$.
+(6)
 where $\ell_{ce}$}} denotes the weighted cross-entropy loss, with weights computed from class frequencies.
-${{\ell_{occ}}}$ is a comprehensive occupancy loss, following prior work [4, 37, 38].
+$\ell_{occ}$}} is a comprehensive occupancy loss, following prior work [22, 17].
 $\bf{\bar{G}}$}} represents the ground-truth of geometry.
-We apply ${{\ell_{occ}}\left({{\bf{S}},{\bf{\bar{S}}}}\right)}$ only to voxels containing more than one point, supervising voxel-level semantic learning.
-3.5.2 Training Strategy
-The training strategy of OccLE is summarized in Algorithm 1.
-In Phases I and II, we independently train the semantic and geometric branches.
-In Phase III, all modules are jointly optimized while freezing the geometric branch to prevent its overfitting.
-Input: Anotated dataset ${{{\cal D}_{l}}}$, unanotated dataset ${{{\cal D}_{u}}}$, Semantic branch ${{\Theta_{sem}}}$, Geometric branch ${{\Theta_{geo}}}$, Semantic-geometric feature grid fusion module ${{\Theta_{fus}}}$
-Output: Trained ${{\Theta_{sem}}}$,${{\Theta_{geo}}}$,${{\Theta_{fus}}}$
-Phase I: Train the semantic branch
-Train ${{\Theta_{sem}}}$ with dataset ${{{\cal D}_{l}}}$,${{{\cal D}_{u}}}$, loss ${{\ell_{sem}}}$, label ${\{{\bf{\bar{S}}},{\bf{\bar{s}}}\}}$
-Phase II: Train the geometric branch
-Train the teacher model ${{\Theta_{geo-T}}}$ with dataset ${{{\cal D}_{l}}}$, loss ${{\ell_{sem}}}$, label ${\{{\bf{\bar{G}}}\}}$
-Predict ${\{{{\bf{\bar{G}}}_{T}}\}}$ for dataset ${{{\cal D}_{u}}}$ via ${{\Theta_{geo-T}}}$
-Train the student model ${{\Theta_{geo-S}}}$ with dataset ${{{\cal D}_{l}}}$,${{{\cal D}_{u}}}$, loss ${{\ell_{sem}}}$, label ${\{{\bf{\bar{G}}}\}}$,${\{{{\bf{\bar{G}}}_{T}}\}}$
-Phase III: Train the semantic-geometric feature grid fusion module
-Freeze ${{\Theta_{geo-S}}}$; Train ${{\Theta_{fus}}}$ and ${{\Theta_{sem}}}$ with dataset ${{{\cal D}_{l}}}$,${{{\cal D}_{u}}}$, loss ${{\ell_{fus}}}$,${{\ell_{sem}}}$, label ${\{{\bf{\bar{Y}}},{\bf{\bar{s}}}\}}$
-Algorithm 1 The training strategy of OccLE
-4 Experiment
-4.1 Dataset and Metric
-4.1.1 Dataset
-We evaluate OccLE on the SemanticKITTI [27], which includes 22 outdoor driving scenarios: 10 sequences for training, 1 for validation, and 11 for testing.
-The dataset annotates scenes of size ${51.2m\times 51.2m\times 6.4m}$ with a voxel size of ${0.2m\times 0.2m\times 0.2m}$ for 20 semantic classes (19 semantics + 1 free).
-To simulate a label-efficient learning, we use only 10% of the voxel annotations from the training set, obtained via interval sampling.
-4.1.2 Metric
-There are two domain metrics: intersection over union (IoU) and mean intersection over union (mIoU) across 19 classes.
-IoU measures the overall scene completion quality, while mIoU evaluates the quality of semantic segmentation for each class.
-We use IoU as the primary metric for the geometric branch and mIoU as the primary metric for the semantic branch and semantic-geometric feature grid fusion.
-4.2 Implementation Details
-[…]
-4.3 Main Result
-We present a quantitative comparison on the SemanticKITTI validation set in Table 2.
-OccLE obtains competitive performance, achieving 16.59 % mIoU and 39.96 % IoU.
-It surpasses all fully supervised camera-based methods (HASSC [40] with 40.66 % mIoU) and approaches the performance of fully supervised multi-modal methods (OccLoff [22] with 22.62 % mIoU).
-As illustrated in Figure 4, OccLE achieves strong performance in complex scenes, accurately reconstructing buildings and vegetation, and distinguishing small objects such as traffic signs and poles.
-Table 2: Quantitative results on SemanticKITTI validation set. Bold and underline represent the best and second best results, respectively.
+We apply ${\ell_{occ}}\left({{\bf{S}},{\bf{\bar{S}}}}\right)$} only to voxels with more than one point to supervise voxel-level semantics.
+Table 2: Quantitative results on SemanticKITTI validation set.
+Bold and underline represent the best and second best results, respectively.
+Inp.
+and Sup.
+indicate the input modality and the supervision type, respectively.
 Method Inp. Sup. IoU mIoU
@@
 OccLoff [52] C+L Full - 22.62 46.44 2.08 3.91 20.38 8.72 3.88 4.35 0.00 66.25 21.07 43.51 0.57 41.23 15.86 41.20 20.06 46.21 27.60 16.47
-OccFusion [5] C+L Full 58.68 21.92 45.62 2.96 3.51 20.05 8.76 3.16 4.37 0.00 65.67 23.08 36.33 0.00 39.09 15.70 45.53 19.37 45.53 27.57 15.21
+OccFusion [30] C+L Full 58.68 21.92 45.62 2.96 3.51 20.05 8.76 3.16 4.37 0.00 65.67 23.08 36.33 0.00 39.09 15.70 40.68 19.37 45.53 27.57 15.21
 OccGen [38] C+L Full 36.87 13.74 26.83 1.60 2.53 15.49 12.83 3.20 3.37 0.00 61.28 20.42 28.30 0.43 14.49 6.94 20.04 3.94 32.44 4.11 2.77
+OccFusion [30] C+L 10% 30.36 6.03 11.24 0.08 0.09 0.64 0.54 0.53 0.09 0.00 39.18 3.23 15.91 0.06 8.42 1.09 13.51 0.28 18.30 1.28 0.07
 RenderOcc [31] C Self - 8.24 14.83 0.42 0.17 2.47 1.78 0.94 3.20 0.00 43.64 12.54 19.10 0.00 11.59 4.71 17.61 1.48 20.01 1.17 0.88
 OccLE(Ours) C+L 10% 40.60 16.59 37.27 2.76 4.39 0.90 10.88 2.52 0.09 0.00 55.79 19.96 28.85 0.14 24.56 12.53 35.19 16.98 35.29 17.16 10.01
+Table 3: Quantitative results on Occ3D-nuScenes validation set. $\dagger$} indicates the use of camera-visible masks during training.
+Method Inp. Sup. mIoU
+other
+barrier
+bicycle
+bus
+car
+cons. veh.
+motorcycle
+pedestrian
+traffic cone
+trailer
+truck
+drive. surf.
+other flat
+sidewalk
+terrain
+manmade
+vegetation
+MonoScene [4] C Full 6.06 1.75 7.23 4.26 4.26 4.93 9.38 3.98 3.98 3.90 4.45 7.17 14.91 6.32 7.92 7.43 1.01 7.43
+OccFormer [53] C Full 21.93 5.95 21.93 5.85 37.83 17.87 40.44 42.43 7.36 23.88 21.81 20.98 22.38 30.70 55.35 28.36 30.70 18.0
+BEVFormer [26] C Full 26.88 5.94 30.29 12.32 34.40 39.17 14.44 16.45 17.22 9.27 13.90 26.36 50.99 30.96 34.66 22.73 6.76 6.97
+TPVFormer [13] C Full 27.83 7.22 38.90 13.67 40.78 45.90 17.23 19.99 18.85 14.30 26.69 34.17 55.65 35.47 37.55 30.70 19.40 16.78
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+SelfOcc [14] C Self 9.30 0.00 0.15 0.66 5.46 12.54 0.00 0.80 2.10 0.00 0.00 8.25 55.49 0.00 26.30 26.54 14.22 5.60
+OccNeRF [49] C Self 9.53 0.00 0.83 0.82 5.13 12.49 3.50 0.23 3.10 1.84 0.52 3.90 52.62 0.00 20.81 24.75 18.45 13.19
+DistillNeRF [39] C Self 8.93 0.03 1.35 2.08 10.21 10.09 2.56 1.98 5.54 4.62 1.43 7.90 43.02 0.00 16.86 15.02 14.06 15.06
+GaussTR [18] C Self 11.70 - 2.09 5.22 14.07 20.43 5.70 7.08 5.12 3.93 0.92 13.36 39.44 - 15.68 22.89 21.17 21.87
+RenderOcc [31] C Self 23.93 5.69 27.56 14.36 19.91 20.56 11.96 12.42 12.14 14.34 20.81 18.94 68.85 33.35 42.01 43.94 17.36 22.61
+OccLE (Ours) C+L 10% 27.53 28.58 0.00 27.33 37.44 8.13 0.00 17.76 0.00 12.56 20.18 71.51 30.97 38.46 38.94 27.90 24.37 83.95
+4 Experiment
+4.1 Dataset and Metric
+4.1.1 Dataset
+We evaluate OccLE on the SemanticKITTI [1] and Occ3D-nuScenes [37].
+SemanticKITTI annotates scenes of size $51.2m\times 51.2m\times 6.4m$} with 0.2m{0.2m} voxel size for 20 classes (19 semantics + 1 free).
+Occ3D-nuScenes annotate the $80.0m\times 80.0m\times 6.4m$} with 0.4m{0.4m} voxel size for 18 classes (17 semantics + 1 free).
+To simulate a label-efficient learning, we use only 10% of the voxel annotations in the training set, obtained via interval sampling.
+4.1.2 Metric
+There are two domain metrics: intersection over union (IoU) and mean intersection over union (mIoU) across 19 classes.
+IoU measures the overall scene completion quality, while mIoU evaluates the quality of semantic segmentation for each class.
+We use IoU as the primary metric for the geometric branch and mIoU as the primary metric for the semantic branch and semantic-geometric feature grid fusion.
+4.2 Implementation Details
+All experiments are conducted on 2 NVIDIA A6000 GPUs with a batch size of 1.
+The AdamW optimizer is used with an initial learning rate of $2{e^{-4}}$} and a weight decay of $e^{-4}$}}, All training durations are set to 40 epochs, while the teacher model of the geometric branch is set to 100 epochs.
+More details are presented in Appendix D.2.
+4.3 Main Result
+We present a quantitative comparison on the SemanticKITTI validation set and Occ3D-nuScenes validation set in Table 2 and Table 3, respectively.
+For the SemanticKITTI dataset, OccLE obtains competitive performance, achieving 16.59 % mIoU and 39.96 % IoU.
+It surpasses all fully supervised camera-based methods (HASSC [40] with 40.66 % mIoU) and approaches the performance of fully supervised multi-modal methods (OccLoff [52] with 22.62 % mIoU).
+For the Occ3D-nuScenes dataset, OccLE is trained using camera-visible masks for 10% of the training samples and a single-frame input.
+The experimental results show that OccLE achieves 27.53 mIoU, outperforming all self-supervised methods (RenderOcc [31] with 23.93% mIoU) and approaching the performance of fully supervised methods.
+As illustrated in Figure 4, OccLE achieves strong performance in complex scenes, accurately reconstructing buildings and vegetation, and distinguishing small objects such as traffic signs and poles.
+Additional comparisons on the SemanticKITTI hidden test set and SSCBenchKITTI-360 [23] validation set are presented in Appendix D.3.
 Figure 4: Qualitative results on the SemanticKITTI validation set.
@@
 4.4 Ablation Studies
-We conduct all ablation studies on the validation set of the SemanticKITTI dataset. We focus on three aspects: semantic branch, geometric branch, and overall pipeline.
-4.4.1 The Semantic Branch
-To evaluate the extractor and pseudo label in the semantic branch, for comparison, we adopt ESPNetv2 [43] and MobileNet [44] as ${{{\cal G}_{s2d}}}$, employ 1-layer and 2-layer versions of sparse UNet3D as ${{{\cal G}_{s3d}}}$, and train using unaligned pseudo labels.
-As depicted in Table 3, 2D extractors designed for edge device suffer a notable performance degradation, whereas our method achieves the highest mIoU(${{\bf{s}}}$) while retaining remaining lightweight.
-Increasing the depth of the sparse UNet3D architecture improves mIoU(${{\bf{S}}}$) without substantially increasing computational cost.
-Moreover, aligned pseudo labels yield a significant mIoU gain by providing supervision for unaligned classes.
+We conduct all ablation studies on the validation set of the SemanticKITTI dataset. Additional ablation studies are presented in the Appendix D.4.
 Table 4: Ablation study of the semantic branch. The unit of Inf. Time is in milliseconds.
@@
 Ours (w/ align.) 6.87 51.51 Ours (w/ align.) 44.59 13.94
+4.4.1 The Semantic Branch
+To evaluate the extractor and pseudo label in the semantic branch, for comparison, we adopt ESPNetv2 [28] and MobileNet [12] as ${\cal G}_{s2d}$}}, employ 1-layer and 2-layer versions of sparse UNet3D as ${\cal G}_{s3d}$}}, and train using unaligned pseudo labels.
+As depicted in Table 4, 2D extractors designed for edge device exhibit a clear drop in performance, whereas our method achieves the highest mIoU($\bf{s}$}}) while remaining lightweight.
+Deeper sparse UNet3D architectures improves mIoU($\bf{S}$}}) without substantially increasing computational cost.
+Aligned pseudo labels significantly improve mIoU by offering supervision for unaligned classes.
 4.4.2 The Geometric Branch
 To evaluate the effects of model architecture and training strategy in the geometric branch, we design a baseline that employs Conv3D and 3D DA for feature extraction and interaction, and compare our semi-supervision against full supervision.
-As depicted in Table 4(a), our cross-plane image and LiDAR feature synergy only has a minor precision drop while significantly reducing module inference time from 138.62 ms to 26.08 ms.
+As depicted in Table 5, our cross-plane image and LiDAR feature synergy only has a minor precision drop while significantly reducing module inference time from 138.62 ms to 26.08 ms.
 Our data augmentation strategy boosts student model performance as augmentation degree increases.
-Under semi-supervision, our student model reaches 53.60 IoU, closely approaching the fully supervised performance of 56.77 IoU.
-Table 4: Ablation studies for (a) geometric branch and (b) overall pipeline.
-“weak,” “medium,” and “strong” denote the degrees of data augmentation, -S and -T represents the student and teacher models.
-‘Sem.’, ‘Geo.’, and ‘Proj.’ denote the semantic branch, geometric branch, and supervising with scatter-accumulated projection, respectively.
-(a) Geometric branch.
-Sup. Weak Middle Strong Model IoU
-Full Ours-T 56.77
-10% Baseline 52.79
-10% Ours-T 52.66
-10% Ours-S 52.86
-10% Ours-S 52.88
-10% Ours-S 53.60
-(b) Overall pipeline.
+Under semi-supervision, our student model reaches 53.60 IoU.
+Table 5: Ablation study of the geometric branch, where “weak,” “medium,” and “strong” denote the degrees of data augmentation, -S and -T represents the student and teacher models.
+Sup. Weak Middle Strong Model Inf. Time IoU
+Full Ours-T 26.08 56.77
+10% Baseline 138.62 52.79
+10% Ours-T 26.08 52.66
+10% Ours-S 26.08 52.86
+10% Ours-S 26.08 52.88
+10% Ours-S 26.08 53.60
+4.4.3 Overall Pipeline
+We use only 10% voxel annotations for supervision as the baseline.
+As depicted in Table 6, our method increases mIoU from 9.64 to 16.59 by incorporating a decoupled design and 2D-aligned pseudo labels, closely approaching the fully supervised performance of 20.32 mIoU.
+Using only aligned pseudo labels significantly reduces precision, indicating pattern collapse when voxel annotations are missing.
+Furthermore, ablating any module in OccLE results in a precision drop, demonstrating the essential role of these three modules in label-efficient learning.
+Table 6: Ablation study of overall pipeline, where ‘Sem.’, ‘Geo.’, and ‘Self.’ denote the semantic branch, geometric branch, and self-supervision in the semantic-geometric fusion module, respectively.
 Sup. Sem. Geo. Proj. Model mIoU IoU
@@
 10% Baseline 9.64 22.30
+10% Ours 0.37 7.84
 10% Ours 11.84 27.46
@@
 10% Ours 16.59 40.60
-4.4.3 Overall Pipeline
-As depicted in Table 4(b), compared to the baseline, our method increases mIoU from 9.64 to 16.59 using only 10% voxel annotations, closely approaching the fully supervised performance of 20.32 mIoU.
-Furthermore, ablating any module in OccLE results in a precision drop, demonstrating the essential role of these three modules in label-efficient learning.
-Additional ablation studies are presented in the Appendix C.2.
+4.4.4 Voxel Annotation Ratio
+As shown in Table 7, OccLE achieves 18.40% mIoU with only 15% voxel annotations, approaching the fully supervised camera+LiDAR method OccLoff [52] (22.62%).
+With just 5% of training data, it still reaches 13.11% mIoU, showing strong performance under limited supervision.
+Table 7: Ablation study of voxel annotation ratio.
+Ratio 5% 10% 15%
+IoU 32.88 40.60 42.10
+mIoU 13.11 16.59 18.40
 5 Conclusion
@@
 OccLE incorporates three key components: (1) distilling 2D foundation model to predict aligned pseudo label for supervising 2D and 3D semantic learning, (2) cross-plane image and LiDAR feature synergy for efficient geometry learning under semi-supervision, and (3) semantic-geometric feature grid fusion with Dual Mamba and scatter-accumulated projection for supervising unannotated regions.
-Experimental results demonstrate that OccLE achieves an mIoU of 16.59% on the SemanticKITTI validation set using only 10% of the voxel annotations.
+Experiments show that OccLE achieves competitive performance with only 10% of voxel annotations on the SemanticKITTI and Occ3D-nuScenes datasets.
 Appendix A Dual Mamba Structure
@@
 We employ dataset-specific and open-vocabulary segmentation 2D foundation models to generate aligned pseudo labels.
-The class mappings are listed in Table B.1.
+The class mappings are listed in Table B.1 and Table B.2.
 We utilize SAM2 [32] to produce trunk class labels and integrate them into the MSeg [19] outputs.
 This example illustrates our aligned pseudo label generation process, which can scale to any dataset with different class taxonomies.
-Table B.1: The aligned label generation.
+Table B.1: The aligned label generation on the SemanticKITTI dataset.
 SemantiKITTI Image Segmentation SemantiKITTI Image Segmentation
@@
 11 sidewalk 100 sidewalk_pavement
-Appendix C Supplementary Experimental Results
-C.1 Supplementary Ablation Study
-We evaluate some experimental setting in Table C.2.
+Table B.2: The aligned label generation on the Occ3D-nuScenes dataset.
+Occ3D-nuScenes Image Segmentation Occ3D-nuScenes Image Segmentation
+Label Class Label Class Label Class Label Class
+0 other rest sky et.al 8 traffic cone SAM2 traffic cone
+1 barrier 130 streetlight 9 trailer 183 trailer
+131 road barrier 10 truck 182 trailer
+144 fence 11 drive. surf. 98 road
+145 railing banister 12 other flat 96 playing field
+145 guard rail 97 railroad
+2 bicycle 175 bicycle 13 sidewalk 100 sidewalk pavemen
+3 bus 180 bus 14 terrain 102 terrain
+4 car 176 car 15 manmade 191 wall
+5 cons. veh SAM2 cons. veh 192 window
+6 motorcycle 178 motorcycle 132-141 traffic sign et al.
+128 motorcyclist 16 vegetation 174 vegetation
+7 pedestrian 125 person
+Appendix C Training Strategy
+The training strategy of OccLE is summarized in Algorithm 1.
+In Phases I and II, we independently train the semantic and geometric branches.
+In Phase III, all modules are jointly optimized while freezing the geometric branch to prevent its overfitting.
+Algorithm 1 The training strategy of OccLE
+0:  Annotated dataset $\mathcal{D}_{l}$, unannotated dataset $\mathcal{D}_{u}$, semantic branch $\Theta_{\text{sem}}$, geometric branch $\Theta_{\text{geo}}$, fusion module $\Theta_{\text{fus}}$
+0:  Trained $\Theta_{\text{sem}},\Theta_{\text{geo}},\Theta_{\text{fus}}$
+1:  Phase I: Train the semantic branch Train $\Theta_{\text{sem}}$ with $\mathcal{D}_{l}$, $\mathcal{D}_{u}$, loss $\ell_{\text{sem}}$, label $\{\bar{\mathbf{S}},\bar{\mathbf{s}}\}$
+2:  Phase II: Train the geometric branch Train teacher model $\Theta_{\text{geo-T}}$ with $\mathcal{D}_{l}$, loss $\ell_{\text{sem}}$, label $\{\bar{\mathbf{G}}\}$ Predict $\{\bar{\mathbf{G}}_{T}\}$ for $\mathcal{D}_{u}$ via $\Theta_{\text{geo-T}}$ Train student model $\Theta_{\text{geo-S}}$ with $\mathcal{D}_{l}$, $\mathcal{D}_{u}$, loss $\ell_{\text{sem}}$, labels $\{\bar{\mathbf{G}}\}$, $\{\bar{\mathbf{G}}_{T}\}$
+3:  Phase III: Train the fusion module Freeze $\Theta_{\text{geo-S}}$ Train $\Theta_{\text{fus}}$ and $\Theta_{\text{sem}}$ with $\mathcal{D}_{l}$, $\mathcal{D}_{u}$, losses $\ell_{\text{fus}}$, $\ell_{\text{sem}}$, labels $\{\bar{\mathbf{Y}},\bar{\mathbf{s}}\}$
+Appendix D Supplementary Experiments
+Table D.3: Quantitative results on SemanticKITTI hidden test set.
+Bold and underline represent the best and second best results, respectively.
+Inp.
+and Sup.
+indicate the input modality and the supervision type, respectively.
+Method Inp. Sup. IoU mIoU
+car (3.92%)
+bicycle (0.03%)
+motorcycle (0.03%)
+truck (0.16%)
+other-veh. (0.20%)
+person (0.07%)
+bicyclist (0.07%)
+motorcyclist (0.05%)
+road (15.30%)
+parking (1.12%)
+sidewalk (11.13%)
+other-grnd.(0.56%)
+building (14.10%)
+fence (3.90%)
+vegetation (39.3%)
+trunk (0.51%)
+terrain (9.17%)
+pole (0.29%)
+traf.-sign (0.08%)
+TPVFormer [13] C Full 34.25 11.26 19.20 1.00 0.50 3.70 2.30 1.10 2.40 0.30 55.10 27.40 27.20 6.50 14.80 11.00 13.90 2.60 20.40 2.90 1.50
+OccFormer [53] C Full 34.53 12.32 21.30 1.50 1.70 3.90 3.20 2.20 1.10 0.20 55.90 31.50 30.30 6.50 15.70 11.90 16.80 3.90 21.30 3.80 3.70
+VoxFormer [22] C Full 43.21 13.41 21.70 1.90 1.60 3.60 4.10 1.60 1.10 0.00 54.10 25.10 26.90 7.30 23.50 13.10 24.40 8.10 24.20 6.60 5.70
+SurroundOcc [42] C Full 34.72 11.86 20.30 1.60 1.20 1.40 4.40 1.40 2.00 0.10 56.90 30.20 28.30 6.80 15.20 11.30 14.90 3.40 19.30 3.90 2.40
+H2GFormer [41] C Full 43.52 14.60 23.70 0.60 1.20 5.20 5.00 1.10 0.10 0.00 57.90 30.00 30.40 6.90 24.00 14.60 25.20 10.70 25.80 7.50 7.10
+SGN [29] C Full 45.42 15.76 25.40 4.50 0.90 4.50 3.70 0.50 0.30 0.10 60.40 28.90 31.40 8.70 28.40 18.10 27.40 12.60 28.40 10.00 8.30
+LowRankOcc [54] C Full 38.47 13.56 20.90 3.30 2.70 2.90 4.40 2.40 1.70 2.30 52.80 25.10 27.20 8.80 22.10 14.40 22.90 8.90 20.80 7.00 7.00
+Symphonies [17] C Full 42.19 15.04 23.60 3.60 2.60 3.20 5.60 3.20 1.90 2.00 58.40 26.90 29.30 11.70 24.70 16.10 24.20 10.00 23.10 7.70 8.00
+HASSC [40] C Full 42.87 14.38 23.00 1.90 1.50 2.90 4.90 1.40 3.00 0.00 55.30 25.90 29.60 11.30 23.10 14.30 24.80 9.80 26.50 7.00 7.10
+Bi-SSC [43] C Full 45.10 16.73 25.00 1.80 2.90 6.80 6.80 1.70 3.30 1.00 63.40 31.70 33.30 11.20 26.60 19.40 26.10 10.50 28.90 9.30 8.40
+OccLE (Ours) C+L 10% 31.42 16.30 29.60 3.50 2.70 0.00 6.30 1.90 0.00 0.00 56.90 22.60 30.80 2.30 24.90 21.90 30.00 19.10 27.90 15.10 14.00
+D.1 Supplementary Datasets
+We evaluate OccLE on additional datasets: SSCBenchKITTI-360[23].
+SSCBenchKITTI-360 shares its scene and voxel configuration with SemanticKITTI and annotates voxels for 19 class labels (18 Semantic + 1 Free).
+D.2 Supplementary Implementation Details
+In the semantic branch, the downsampling scale of ${\cal G}_{s2d}$}} and ${\cal G}_{s3d}$}} is $8\times$}; In the geometric branch, we utilize ResNet50 [10] and a 2-layer sparse Conv3D to extract features from image and LiDAR scans.
+In the semantic-geometric feature grid fusion module, we stack 4 Dual Mamba blocks.
+For the SemanticKITTI dataset, we adopt the multi-frame setting as in [22, 29, 41, 40], and crop the camera images to $1220\times 370$}.
+For the Occ3D-nuScenes dataset, we use a single-frame setting and crop the images to $1600\times 900$}.
+For the SSCBenchKITTI-360 dataset, we follow the same multi-frame setting as SemanticKITTI and crop the images to $1408\times 376$}.
+D.3 Supplementary Quantitative Comparison
+The quantitative comparison results on the SemanticKITTI hidden test set and SSCBenchKITTI-360 validation set are presented in Table D.3 and Table D.4.
+OccLE achieves 16.30 mIoU and 31.42 IoU on the SemanticKITTI hidden test set, ranking second among camera-based fully supervised methods.
+On the SSCBenchKITTI-360 validation set, OccLE achieves 16.38 mIoU and 52.44 IoU, performing competitively against both camera-based and LiDAR-based methods.
+These results indicate that OccLE can perform well even with limited voxel-level annotations.
+Table D.4: Quantitative results on SSCBench-KITTI-360 validation set.
+Method Inp. Sup. IoU mIoU
+car (2.85%)
+bicycle (0.01%)
+motorcycle (0.01%)
+truck (0.16%)
+other-veh. (5.75%)
+person (0.02%)
+road (14.98%)
+parking (2.31%)
+sidewalk (6.43%)
+other-grnd.(2.05%)
+building (15.67%)
+fence (0.96%)
+vegetation (41.99%)
+terrain (7.10%)
+pole (0.22%)
+traf.-sign (0.06%)
+other-struct. (4.33%)
+other-obj. (0.28%)
+TPVFormer [13] C Full 40.22 13.64 21.56 1.09 1.37 8.06 2.57 2.38 52.99 11.99 31.07 3.78 34.83 4.80 30.08 17.52 7.46 5.86 5.48 2.70
+OccFormer [53] C Full 40.27 13.81 22.58 0.66 0.26 9.69 3.82 2.77 54.30 13.44 31.53 3.55 36.42 4.80 31.00 19.51 7.77 8.51 6.95 4.60
+VoxFormer [22] C Full 38.76 11.91 17.84 1.16 0.89 4.56 2.06 1.63 47.01 9.67 27.21 2.89 31.18 4.97 28.99 14.69 6.51 6.92 3.79 2.43
+SGN [29] C Full 47.06 18.25 29.03 3.43 2.90 10.89 5.20 2.99 58.14 15.04 36.40 4.43 42.02 7.72 38.17 23.22 16.73 16.38 9.93 5.86
+Symphonies [17] C Full 44.12 18.58 30.02 1.85 5.90 25.07 12.06 8.20 54.94 13.83 32.76 6.93 35.11 8.58 38.33 11.52 14.01 9.57 14.44 11.28
+GaussianFormer [16] C Full 35.38 12.92 18.93 1.02 4.62 18.07 7.59 3.36 45.47 10.89 25.03 5.32 28.44 5.68 29.54 8.62 2.99 2.32 9.51 5.14
+GaussianForme-2 [15] C Full 38.37 13.90 21.08 2.55 4.21 12.41 5.73 1.59 54.12 11.04 32.31 3.34 32.01 4.98 28.94 17.33 3.57 5.48 5.88 3.54
+LMSCNet[33] L Full 47.53 13.65 20.91 0.00 0.00 0.26 0.00 0.00 62.95 13.51 33.51 0.20 43.67 0.33 40.01 26.80 0.00 0.00 3.63 0.00
+SSCNet[36] L Full 53.58 16.95 31.95 0.00 0.17 10.29 0.58 0.07 65.70 17.33 41.24 3.22 44.41 6.77 43.72 28.87 0.78 0.75 8.60 0.67
+OccLE (Ours) C+L 10% 52.44 16.38 31.14 0.34 0.28 4.66 1.39 2.05 53.47 9.15 29.65 4.32 45.07 7.30 41.53 23.79 19.81 11.36 6.45 3.11
+D.4 Supplementary Ablation Study
+We evaluate some experimental setting in Table D.2.
 To assess the impact of semi-supervision in the geometric branch on the overall pipeline, we incorporated a teacher model of the geometric branch into final joint training.
@@
 This result demonstrates that the scatter-accumulated projection is simple yet effective in supervising all voxels with aligned pseudo labels, whereas the distance-weighted projection reduces supervision quality for distant voxels.
-We assess the necessity of using limited voxel annotations rather than relying solely on aligned pseudo labels.
-Our experiments indicate that relying solely on aligned pseudo labels results in a marked decline in precision, signaling pattern collapse during training when voxel annotations are absent.
-These results confirm that voxel annotations are essential for guiding the model toward an effective fusion of semantic and geometric features.
-Table C.2: Supplementary ablation study.
+Table D.2: Ablation study of pipeline and component.
 Model mIoU IoU
@@
 w/ weighted projection 13.38 40.40
-w/o limited voxel annotation 0.37 7.84
 Ours 16.59 40.60
-C.2 Supplementary Qualitative Comparison
+D.5 Supplementary Qualitative Comparison
 We show additional qualitative results of OccLE on the SemanticKITTI validation dataset in Figure C1.
@@
 Figure C1: The qualitative results of OccLE on SemanticKITTI validation dataset.
-Appendix D Limitation
+Appendix E Limitation
 In this study, we simulate label-efficient learning for 3D semantic occupancy prediction by uniformly sampling 10% of voxel annotations across all scenes in the SemanticKITTI dataset.